### v1

# Temporal Action Localization for Inertial-based Human Activity Recognition

Abstract.
A persistent trend in Deep Learning has been the applicability of machine learning concepts to other areas than originally introduced for. As of today, state-of-the-art activity recognition from wearable sensors relies on classifiers being trained on fixed windows of data. Contrarily, video-based Human Activity Recognition has followed a segment-based prediction approach, localizing activity occurrences from start to end. This paper is the first to systematically demonstrate the applicability of state-of-the-art TAL models for wearable Human Activity Recongition (HAR) using raw inertial data as input. Our results show that state-of-the-art TAL models are able to outperform popular inertial models on 4 out of 6 wearable activity recognition benchmark datasets, with improvements ranging as much as $25\%$ in F1-score. Introducing the TAL community’s most popular metric to inertial-based HAR, namely mean Average Precision, our analysis shows that TAL models are able to produce more coherent segments along with an overall higher NULL-class accuracy across all datasets. Being the first to provide such an analysis, the TAL community offers an interesting new perspective to inertial-based HAR with yet to be explored design choices and training concepts, which could be of significant value for the inertial-based HAR community.
Deep Learning, Inertial-based Human Activity Recognition, Body-worn Sensors, Temporal Action Localization
ccs: Human-centered computing Ubiquitous and mobile computing design and evaluation methodsccs: Computing methodologies Neural networks
1. Introduction
The recognition of performed activities through wearable sensors such as IMUs has shown to be of significant value in areas such as health care or the support of complex work processes (Bulling et al., 2014). Deep Learning has established itself as the de facto standard in wearable, inertial-based Human Activity Recognition (HAR), consistently outperforming classical Machine Learning approaches in recognition performance. A persistent trend in Deep Learning has been the applicability of machine learning concepts such as self-attention (Vaswani et al., 2017) to other areas and application scenarios than originally introduced for. With significant progress having been made since the introduction of deep neural architectures such as the DeepConvLSTM (Ordóñez and Roggen, 2016), researchers have followed this trend and continuously worked on improving the architectural design of networks by incorporating newly introduced techniques (see e.g., (Zhou et al., 2022)).
A promising recent approach in video-based Human Activity Recognition (HAR) is Temporal Action Localization (TAL), which aims to locate activity segments, defined by a class label, start, and end point, within an untrimmed video. Even though introduced architectures have almost doubled in performance over the last 5 years on existing datasets like THUMOS-14 (Jiang et al., 2014), results on large corpora such as EPIC-KITCHENS-100 (Damen et al., 2022) and Ego4D (Grauman et al., 2022) show that the prediction problem is far from being saturated. Recently, Zhang et al. (2022b) introduced the ActionFormer, an end-to-end, single-stage TAL model utilizing transformer-based layers. Being the first of its kind, the ActionFormer along with extensions of it (Tang et al., 2023; Shi et al., 2023) are as of today holding top ranks in prediction performance on a multitude of TAL benchmark datasets. Despite sharing a mutual goal, the TAL community evaluates their approaches differently from the inertial community, using mean Average Precision (mAP) as opposed to traditional classification metrics like F1-score or accuracy. TAL models have recently been shown to be capable of being trained using raw inertial data (Bock et al., 2023), marking the first instance of such vision-based models being applied in the context of inertial-based HAR. When compared with popular inertial models, results on one particular dataset have already shown that TAL models can produce more coherent and less fragmented predictions while maintaining performance in terms of traditional classification metrics.
Setting out to further examine the applicability of Temporal Action Localization for inertial-based HAR, our contributions in this paper are three-fold:
  1. (1)
    We demonstrate the capabilities of a novel approach to inertial-based HAR, being single-stage TAL, to outperform popular inertial models on four out of six wearable activity recognition benchmark datasets.
  2. (2)
    Complementing traditional, frame-wise classification metrics, we introduce a set of unexplored, segment-based evaluation metrics for inertial-based HAR, which are based around the scalar evaluation metric mean Average Precision (Alwassel et al., 2018).
  3. (3)
    We name differences and similarities between models originating from both communities, highlighting applied concepts that TAL models use to produce coherent segments along with a high NULL-class accuracy.
2. Related Work
Inertial-based Human Activity Recognition
With on-body sensors providing a robust and non-intrusive way to monitor participants along long stretches of time, research conducted in the area of inertial-based HAR has worked towards the automatic interpretation of one or multiple sensor streams to reliably detect activities e.g. in the context of providing medical support or providing guidance during complex work processes (Bulling et al., 2014). With deep neural networks having established themselves as the de facto standard in inertial-based HAR, the DeepConvLSTM (Ordóñez and Roggen, 2016) as well as models introduced at a later point all follow a similar prediction scenario design applying a sliding window approach which groups concurrent data points for classification. In their original publication Ordóñez and Roggen (2016) have found a combination of convolutional and recurrent layers to produce competitive results, with the idea being to model temporal dependencies amongst automatically extracted discriminative features within a sliding window in order to classify it correctly as one of the $N$ activity classes or, if applicable, NULL-class. Building up on the idea of combining these two types of layers, researchers have worked on extending the original DeepConvLSTM or have introduced their own architectural designs (Murahari and Plötz, 2018; Vaizman et al., 2018; Peng et al., 2018; Xi et al., 2018; Xu et al., 2019; Xia et al., 2020; Liu et al., 2020; Chen et al., 2020; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022; Zhang et al., 2022a; Miao et al., 2022; Chen et al., 2023b, a; Wieland and Pankratius, 2023). A simple modification of the DeepConvLSTM is the shallow DeepConvLSTM (Bock et al., 2021). Contradicting the popular belief that one needs at least two recurrent layers when dealing with time series data (Karpathy et al., 2015), Bock et al. (2021) demonstrated that removing the second LSTM layer within the original DeepConvLSTM architecture, results in significant increases in predictive performance on a multitude of HAR benchmark datasets while also decreasing the number of learnable parameters. Furthermore, with the original DeepConvLSTM only being able to learn temporal dependencies within a sliding window, the shallow DeepConvLSTM applies the LSTM across batches, effectively making the batches the sequence which is to be learned by the LSTM. This dimension-flip, along with a non-shuffled training dataset, enables the architecture to learn temporal dependencies amongst a batched input. The same year as the shallow DeepConvLSTM, Abedin et al. (2021) introduced Attend-and-Discriminate, a deep neural network architecture following the idea of the original DeepConvLSTM by combining both convolutional and recurrent layers, yet further adding a cross-channel interaction encoder using self-attention as well as attention mechanism to the recurrent parts of the network. In 2022 Zhou et al. (2022) proposed TinyHAR, a lightweight HAR model that uses a transformer encoder block along with means of cross-channel fusion via a fully connected layer and a final self-attention layer which aims to learn the contribution of each outputted time step produced by the recurrent parts individually.
Video-based Human Activity Recognition
Classifying videos in the context of Human Activity Recognition can broadly be categorized into three main application scenarios: Action Recognition, which aims to classify trimmed videos into one of $C$ activity classes; Action Anticipation, which aims to predict the next likely activities after observing a preceding video sequence; and Temporal Action Localization (TAL), which seeks to identify and locate all activity segments within an untrimmed video. With the inertial-based benchmark datasets consisting of a multitude of activities, TAL is to be considered most comparable to inertial-based HAR. Unlike popular intertial architectures though, TAL models aim to predict all segments within an untrimmed video at once. Existing TAL methods can broadly be categorized into two categories: two-stage and one-stage recognition systems. Two-stage recognition system (Lin et al., 2019, 2020; Xu et al., 2020; Bai et al., 2020; Zhao et al., 2020; Zeng et al., 2019; Gong et al., 2020; Liu et al., 2021; Qing et al., 2021; Sridhar et al., 2021; Zhu et al., 2021; Zhao et al., 2021; Tan et al., 2021) divide the process of identifying actions within an untrimmed video into two subtasks. First, during proposal generation, candidate segments are generated, which are then, during the second step, classified into one of $C$ activity classes and iteratively refined regarding their start and end times. Contrarily, single-stage models (Yang et al., 2023; Shi et al., 2022; Nag et al., 2022; Liu et al., 2022; Liu and Wang, 2020; Long et al., 2019; Lin et al., 2021; Chen et al., 2022; Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) do not rely on activity proposals, directly aiming to localize segments along their class label and refined boundaries. In 2022 Zhang et al. (2022b) introduced the ActionFormer, a lightweight, single-stage TAL model which unlike previously introduced single-stage architectures does not rely on pre-defined anchor windows. In line with the success of transformers in other research fields, Zhang et al. (2022b) demonstrated their applicability for TAL, outperforming previously held benchmarks on several TAL datasets (Heilbron et al., 2015; Damen et al., 2022; Jiang et al., 2014) by a significant margin. Surprisingly, a year later the TemporalMaxer suggested removing transformer layers within the ActionFormer arguing feature embeddings are already discriminative enough (Tang et al., 2023). Though being more lightweight than the ActionFormer, the TemporalMaxer showed to outperform its precedent during benchmark analysis. Similar to the TemporalMaxer, Shi et al. (2023) introduced TriDet, which suggested altering the ActionFormer in two ways. First, to mitigate the risk of rank-loss, self-attention layers are replaced with SGP layers. Second, the regression head in the decoder is replaced with a Trident head, which improves imprecise boundary predictions via an estimated relative probability distribution around each timestamp’s activity boundaries.
3. Temporal Action Localization for Inertial-based HAR
As the inertial-based HAR and TAL communities deal with inherently different modalities, both communities have developed distinct predictive pipelines and algorithms tailored to the challenges and characteristics of their respective modalities (see Figure 1). In the following section, we will provide detailed descriptions of the two different predictive pipelines and demonstrate how vision-based TAL models can be applied to a previously unexplored modality — inertial data. Furthermore, we provide a detailed overview of the architectural design of three popular single-stage TAL models (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) highlighting similarities as well as differences of the architectures compared to models originating from the inertial-based HAR community.
3.1. Inertial-based HAR vs. Temporal Action Localization
Figure 1. Overview of the prediction pipelines applied in inertial-based activity recognition and single-stage Temporal Action Localization (TAL). Both apply a sliding window to divide input data into windows of a certain duration (e.g. one second). TAL models do not use raw data as input but are applied on per-clip, pre-extracted feature embeddings. Inertial activity recognition models predict activity labels for each sliding window, which are used for calculating classification metrics such as accuracy and F1. TAL models predict activity segments, defined by a label, start and end points, and are evaluated with mean Average Precision (mAP) applied at different temporal Intersection over Union (tIoU) thresholds.
Problem definition
The objective of both inertial activity recognition and TAL is to predict all activities present in an untrimmed timeline. Given an input data stream $X$ of a sample participant, both the inertial and TAL communities start by applying a sliding window approach which shifts over $\mathbf{X}$, dividing the input data into windows, e.g. of one second duration with a 50% overlap between consecutive windows. This process results in $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}$ being discretized into $t=\{0,1,...,T\}$ time steps, where $T$ is the number of windows for each participant. It is important to note that the TAL models do not use raw data as input, but are instead trained using feature embeddings extracted from each individual sliding video clip, which are extracted using pretrained methods such as (Carreira and Zisserman, 2017). That is, in the case of TAL, $\mathbf{x}_{t}$ represents a one-dimensional feature embedding vector associated with the video clip at time step $t$. In case of inertial-based activity recognition, $\mathbf{x}_{t}$ represents a two-dimensional matrix of size $\mathbb{R}^{W\times S}$, where $W$ is the number of samples within a window, and $S$ is the number of sensor axes present in the sensor data.
Given all sliding windows associated with an untrimmed sequence, inertial activity recognition models aim to predict an activity label $a_{t}$ for each sliding window $\mathbf{x}_{t}$, where $a$ belongs to a predefined set of activity labels, $a_{t}\in{1,...,C}$. To do so, the sliding windows are batched and fed through e.g. a deep neural network, such as the DeepConvLSTM (Ordóñez and Roggen, 2016). The resulting activity labels for each window are then compared to the true activity labels from the ground truth data, and classification metrics like accuracy or F1-score are calculated. Contrarily, TAL models aim to identify and localize segments of actions within the untrimmed data stream, which can span across multiple windows. To achieve this algorithms are trained to predict activity segments $Y=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N}\}$ using a collection of feature embeddings $X$ as input, where $N$ varies across participants. Each activity instance $\mathbf{y_{i}}=(s_{i},e_{i},a_{i})$ is defined by its starting time $s_{i}$ (onset), ending time $e_{i}$ (offset) and its associated activity label $a_{i}$, where $s_{i}\in[0,T]$, $e_{i}\in[0,T]$ and $a_{i}\in\{1,...,C\}$.
Evaluation metrics
As predictive outcomes of inertial activity recognition and TAL differ, both communities have adopted different evaluation metrics. In case of the inertial community, the problem of recognizing activities is translated to a per-window classification problem, where each window represents an instance to be classified into one of $C$ classes. Therefore, most researchers assess their algorithms based on traditional classification measures, including recall ($R$), precision ($P$), and F1-score ($F1$). In order to calculate these activity-wise metrics one first needs to calculate activity-wise True Positives (TP), False Positives (FP), True Negatives (TN) and False Negatives (FN). For a given activity class $i\in\{1,...,C\}$, a predicted activity label is to be treated as:
  • TP of class $i$ ($TP_{i}$): the predicted and ground truth label are equal to $i$
  • FP of class $i$ ($FP_{i}$): the predicted label is equal to $i$, yet the ground truth is not
  • TN of class $i$ ($TN_{i}$): the predicted and ground truth label are not equal to $i$
  • FN of class $i$ ($FN_{i}$): the predicted label is not equal to $i$, yet the ground truth is
Calculating TP, FP, TN and FN counts of all activities then allows to calculate the per-class classification metrics, which are defined as:
(1) $P_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}}$
(2) $R_{i}=\frac{TP_{i}}{TP_{i}+FN_{i}}$
(3) $F1_{i}=\frac{2\cdot P_{i}\times R_{i}}{P_{i}+R_{i}}$
To account for the continuous nature of activity recognition prediction streams, Ward et al. (2011) introduced additional performance metrics to explicitly measure misalignment within predictions made by inertial-based activity recognition systems. According to the authors, calculating classification metrics such as precision and recall based on window-wise errors does not provide intuitive insights about the overall fragmentation and merging behavior of the prediction stream. The authors propose to further divide the class-wise counts of FPs and FNs into additional sub-categories based on their relative position within a segmented predictive stream. These misalignment sub-categories are:
  1. (1)
    Underfill ($U_{i}$): False Negatives of class $i$ that occur at the start or end of an activity segment, effectively making the predicted segment shorter by starting later or finishing earlier than the corresponding ground truth segment of $i$.
  2. (2)
    Deletion ($D_{i}$): False Negatives of class $i$ that correspond to a deleted event, i.e. a ground segment which is missed completely within the prediction stream.
  3. (3)
    Fragmenting ($F_{i}$): False Negatives of class $i$ that occur in-between True Positive instances of class $i$, falsely splitting a coherent segment into two segments.
  4. (4)
    Insertion ($I_{i}$): False Positives of class $i$ that correspond to a non-existent ground truth segment.
  5. (5)
    Overfill ($O_{i}$): False Positives of class $i$ that occur at the start or end of an activity segment, effectively making the predicted segment longer by predicting an earlier start or later finish than the corresponding ground truth segment.
  6. (6)
    Merge ($M_{i}$): False Positives of class $i$ that occur between two ground truth segments incorrectly merging the two segments into one coherent segment.
Calculating these six misalignment $U$, $D$, $F$, $I$, $O$ and $M$ counts of all activities then allows calculating the per-class misalignment ratios, which are defined as:
(4) $UR_{i}=\frac{U_{i}}{TN_{i}+FN_{i}}$
(5) $DR_{i}=\frac{D_{i}}{TN_{i}+FN_{i}}$
(6) $FR_{i}=\frac{F_{i}}{TN_{i}+FN_{i}}$
(7) $IR_{i}=\frac{I_{i}}{TP_{i}+FP_{i}}$
(8) $OR_{i}=\frac{O_{i}}{TP_{i}+FP_{i}}$
(9) $MR_{i}=\frac{M_{i}}{TP_{i}+FP_{i}}$
Inspired by measures applied during object detection, the TAL community assess algorithms based mean Average Precision (mAP) performance applied at different temporal Intersection over Union (tIoU) thresholds. With a TAL model being trained to predict a set of activity segments $\mathbf{Y}$, each predicted segment is compared with the ground truth activity segments within the complete input data stream of a participant. That is, given a predicted activity segment $\mathbf{y_{p}}=(s_{p},e_{p},a_{p})$ and a ground truth activity segment $\mathbf{y_{g}}=(s_{g},e_{g},a_{g})$, the two segments are considered to be either TP, FP, TN or FN based on the condition that the temporal intersection over union (tIoU) (overlap of the two segments divided by the union of the two segments) is greater than a predefined threshold, e.g., 0.5. Mean Average Precision is then calculated as:
(10) $mAP=\frac{1}{N}\sum^{C}_{a=1}{AP_{a}},$
where $AP_{i}$ represents the area under the curve of the Precision-Recall-Curve, which is obtained by plotting a model’s precision and recall values as a function of the model’s confidence score threshold for a given class $a\in\{1,...,C\}$. More specifically, the mAP score at a tIoU threshold of, for example, 0.2 (mAP@0.2), is calculated as the mAP score assuming a required tIoU of 0.2 when calculating $TP$, $FP$, $TN$ and $TN$ counts. Typically, to assess the predictive performance of TAL models, mAP is calculated using a range of tIoU thresholds, which are then averaged. In case of this benchmark analysis we calculate mAP as mAP@(0.3:0.1:07), which equates to calculating the mAP@0.3, mAP@0.4, mAP@0.5, mAP@0.6 and mAP@0.7 and averaging the individual scores.
3.2. Vectorization of Inertial Data for TAL
As previously mentioned, the input data $X$ used to train TAL models is a collection of feature embeddings. These embeddings are, for instance, extracted using some pre-trained method applied on each video clip. The resulting embeddings are 1-dimensional feature vectors that summarize information such as RGB and optical flow of the video clip (see, for example, Carreira and Zisserman (2017)). Since both communities employ a sliding window approach but feed data to their models using different dimensions, Bock et al. (2023) proposed a simple, yet effective preprocessing step. This step converts the 2-dimensional, windowed inertial data, as used by inertial architectures, into a 1-dimensional feature embedding suitable for training TAL models. The preprocessing method as detailed in Figure 2 involves concatenating the different sensor axes of each window, converting the input data to be a collection of 1-dimensional feature embedding vectors $\mathbf{x}_{t}\in\mathbb{R}^{1\times WS}$ where $W$ is the number of samples within a window, and $S$ is the number of sensor axes. More specifically, given a sliding window $\mathbf{x}_{SW}\in\mathbb{R}^{W\times S}$, we vectorize the two-dimensional matrix as follows:
(11) $\mathbf{x}_{SW}=\begin{bmatrix}x_{11}&\dots&x_{1S}\\ \vdots&\ddots&\\ x_{W1}&&x_{WS}\end{bmatrix}\to\text{vec}(\mathbf{x}_{SW})=\begin{bmatrix}x_{11}\\ \vdots\\ x_{1S}\\ x_{2S}\\ \vdots\\ x_{WS}\end{bmatrix}$
Figure 2. Visualization of the applied vectorization on top of windowed inertial data assuming four 3D-inertial sensors and a sliding window size of 50 samples. Each 2D-sliding-window of size $[50\times 12]$ is vectorized by concatenating each of the axes one after another. Resulting 1D-embedding vectors, being of size $[1\times 600]$, can be used to train TAL models.
3.3. Architectures Overview
In order to predict activity segments $Y=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N}\}$ within an input video, the ActionFormer, TemporalMaxer and TriDet model all follow the same sequence labeling problem formulation for action localization. That is, given a set of feature input vectors $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}$, a model aims to classify each timestamp’s as either one of the activity categories $C$ or as background (NULL-class) and regress the distance towards the timestamp’s corresponding segment’s start and end point. More specifically, given the input vectors $X$ a model aims to learn to label the sequence as
(12) $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}\to\mathbf{\hat{Y}}=\{\mathbf{\hat{y}_{1}},\mathbf{\hat{y}_{2}},...,\mathbf{\hat{y}_{T}}\}$
, where $\mathbf{\hat{y}_{t}}=(p(a_{t}),d_{t}^{s},d_{t}^{e})$ at timestamp $t$ is defined by a probability vector $p(a_{t})$ indicating the class-wise probability of the timestamp being classified as one of the activity categories $C$ and $d_{t}^{s}>0$ and $d_{t}^{e}>0$ being the distance between the current timestamp $t$ and the current segment’s onset and offset. Note that $d_{t}^{s}$ and $d_{t}^{e}$ are not defined if the timestamp is to be classified as background. The sequence labeling formulation can then be easily decoded to activity segments as:
(13) $a_{t}=\arg\max p(a_{t}),\quad s_{t}=t-d_{t}^{s},\quad\textrm{and}\quad e_{t}=t+d_{t}^{e}$
ActionFormer
Introduced in 2022, the ActionFormer was among the first single-stage, anchor-free TAL models which followed a transformer-based design (Zhang et al., 2022b). The architecture, depicted in Figure 3, follows an encoder-decoder structure that decomposes the previously defined task of learning to predict activity segments using a function $f$ as $h\circ g$, where $g:X\to Z$ serves as the encoder and $h:Z\to\hat{Y}$ is the decoder. As the encoder Zhang et al. (2022b) employ a combination of convolutional layers and a transformer network. The convolutional layers, paired with ReLU activation functions, project each input feature vector $\mathbf{x}_{t}$ into a $D$-dimensional space. To limit complexity, each layer of the transformer network applies alternating layers of multi-head local self-attention and MLP. To capture information at different temporal scales, the transformer network is structured in such a way that across its $L$ layers, the initially embedded feature vector $\mathbf{z}_{0}$ at timestamp $t$ is downsampled using a strided depthwise 1D-convolution.
After encoding the input sequence into a feature pyramid $Z=\{\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{L}\}$, which captures information at various temporal scales, the decoder decodes the feature pyramid $Z$ into sequence labels $\hat{Y}$. To do so, the decoder includes both a classification and regression head. The classification head, being 1D convolutional network complemented with layer normalization and ReLU activation, predicts for all $L$ pyramid levels the probability vector $p(a_{t})$ at each timestep $t$. The regression head, following the same design, predicts the onset ($d_{t}^{s}$) and offset distance ($d_{t}^{e}$) at all pyramid levels $L$ for each timestamp $t$. Note that for both the classification and regression heads, parameters are shared across pyramid levels and output regression ranges are pre-specified per pyramid level. Finally, to reduce overlapping instances and create a final set of predicted activity segments, the obtained predictions across all pyramid levels are postprocessed using Soft-NMS (Bodla et al., 2017).
Figure 3. Architecture overview of the ActionFormer proposed by Zhang et al. (2022b). The architecture follows a encoder-decoder structure. The encoder, comprising of a projection using convolutional layers and a transformer network, encodes input sequences into a feature pyramid, which captures information at various temporal scales. The decoder, consisting of a classification and regression head, then decodes each timestamp within the feature pyramid to sequence labels, i.e. a class probability vector and the timestamp’s activity onset and offset distance. Note that weights of the decoder are shared across pyramid levels.
TemporalMaxer
Building upon the architectural design of the Actionformer (Zhang et al., 2022b), the TemporalMaxer proposed by Tang et al. (2023) suggests modifying the encoder of the ActionFormer to employ solely max pooling and remove any multi-head attention layers (see Figure 4). According to the authors, the removal of the transformer-based layers does not come at the cost of a lost in information, as the extracted feature embeddings $X$ outputted by the projection layer are already discriminative enough to achieve state-of-the-art predictive performance. Same as the ActionFormer, the weights of both the classification and regression heads within the TemporalMaxer are shared across pyramid levels, output regression ranges are pre-specified per pyramid level, and to create a final set of predicted activity segments, obtained predictions across all pyramid levels are postprocessed using Soft-NMS (Bodla et al., 2017).
Figure 4. Architecture overview of the TemporalMaxer (Tang et al., 2023). The authors propose to modify the transformer of the ActionFormer (Zhang et al., 2022b) to employ solely max pooling, removing any self-attention mechanisms from the architecture. Each resulting embedded timestamp within the feature pyramid is then decoded using a classification and a regression head. Note that weights of the decoder are shared across pyramid levels.
TriDet
Similar to the TemporalMaxer, the TriDet model (see Figure 5) suggested to make modifications to the architectural design of the ActionFormer by altering both the encoder and regression head decoder (Shi et al., 2023). The authors claim that a transformer-based encoder is prone to suffer from rank loss while simultaneously being high in computational complexity. Consequently, Shi et al. (2023) propose replacing the projection and transformer network of the ActionFormer with Scalable-Granularity Perception (SGP) layers. These layers, being fully-convolutional, comprise of two branches: an instant-level branch, which increases the discriminability between the activity classes and the NULL-class, and a window-level branch, aiming to capture semantic context from a wider receptive field. Mathematically, the SGP operator replacing the mulit-head attention operator is defined as:
(14) $f_{SGP}=\phi(x)FC(x)+\psi(x)(Conv_{w}(x)+Conv_{kw}(x))+x$
, where $FC$ are fully-connected, $Conv_{w}$ are 1D-depth-wise convolutional layers over the window size $w$ and $k$ is a scalable factor controlling the granularity of captured temporal information. In the TriDet model, the $L$ layers of the feature pyramid $Z=\{\mathbf{z}_{1},\mathbf{z}_{2},...,\mathbf{z}_{L}\}$ are thus defined as first downsampling the input feature embeddings $L$ times and applying a SGP layer to each downsampled embedding. According to the authors, the ActionFormer architecture suffers from imprecise boundary predictions. To address this, the original regression head is replaced by a Trident-head, which aims to overcome the ActionFormer’s shortcomings via an estimated relative probability distribution around each activity boundary. The Trident-head consists of three main components: a start head, responsible for locating the start of an activity segment; an end head, responsible locating the end of an activity segment; and a center-offset head, responsible for locating the center of an activity segment. Using these three heads, the Trident-head predicts the start ($d_{t}^{s}$) and end boundary ($d_{t}^{e}$) for all $L$ pyramid levels for each timestamp $t$. Note that all heads, including the classification head, are modelled as three 1D-convolutional layers with parameters shared across all feature pyramid levels.
Figure 5. Architecture overview of the TriDet model proposed by Shi et al. (2023). Compared to the Actionformer (Zhang et al., 2022b), the architecture proposes the following changes: 1. replace the projection and transformer network of the ActionFormer with Scalable-Granularity Perception (SGP) layers along with preceding downsample operators via max-pooling, and 2. replace the regression head with a Trident head, consisting of a start, end and center-offset head. Each resulting embedded timestamp within the feature pyramid is then decoded using the newly defined decoder. Note that weights of the decoder are shared across pyramid levels.
3.4. Similarities and Differences between Inertial and TAL Models
Similar to inertial models, the ActionFormer, TemporalMaxer and TriDet model aim to learn temporal dependencies by learning discriminative feature embeddings. Nevertheless, do the authors accredit much of their models’ performance to the constructed multi-layer feature pyramid. As described above, the feature pyramid within each of the three models downsamples the input embedding multiple times in order to create embedding vectors of different temporal granularity. Though this is similar to downsampling performed by concatenating multiple convolution layers without padding, as demonstrated by all inertial architectures in this benchmark analysis, the TAL models use each embedding vector separately instead of only using the resulting embedding vector of the last pyramid layer. More specifically, the authors claim that predicting timestamps at different degrees of temporal granularity holds significant value and is essential in order to correctly predict different types and lengths of activities. Furthermore, as already detailed in previous chapters, the TAL community aims to localize segments rather than provide window-wise predictions. Due to their objective thus differing from that of the inertial community, TAL models are trained not only on a classification loss, but also a regression loss, which optimizes each timestamps corresponding activity onset and offset. Lastly, the use of self-attention based layers has shown to improve results in both communities, yet variations of the ActionFormer show that these layers are not necessarily needed. More specifically, the use of fully-convolutional SGP layers in the TriDet model show that, rather than the type of technique being important, focus needs to be put specifically learning both local and global temporal information for each timestamp.
4. Methodology
4.1. Datasets
We evaluate each algorithm featured in this benchmark analysis using 6 popular HAR datasets, namely the Opportunity (Roggen et al., 2010), SBHAR (Reyes-Ortiz et al., 2016), Wetlab (Scholl et al., 2015), WEAR (Bock et al., 2023), Hang-Time (Hoelzemann et al., 2023) and RWHAR dataset (Sztyler and Stuckenschmidt, 2016). The datasets, all covering different application and recording scenarios, provide us with a challenging set of prediction problems to properly assess the strengths and weaknesses of each model. Table 1 summarizes the key characteristics of each dataset. In addition to vital information such as participant count, activity count, and sensor axes count, the table also includes details on the type of activities found in each dataset. We classify activities into four types: (1) periodic activities, characterized by recurring periodic patterns; (2) non-periodical activities consisting of non-occurring, non-periodical patterns; and (3) complex activities, defined by an arbitrary sequence of non-periodic and periodic activities. Lastly, following the works of Alwassel et al. (2018), we provide average count of segments across all subjects, categorizing each segment within each dataset into five bins: XS: (0 seconds, 3 seconds], S: (3 seconds, 6 seconds], M: (6 seconds, 12 seconds], L: (12 seconds, 18 seconds], and XL: more than 18 seconds.
Table 1. Key characteristics of the datasets used in this benchmark analysis. The table provides: participant count (#Sbjs), activity count (#Cls), sensor axes count (#Axes), average number of segments per participant based on absolute length of the segment (#Segments) and type of activities found in the dataset.
Dataset #Sbjs #Cls #Axes #Segments (Alwassel et al., 2018) Type of Activities
XS S M L XL
Opportunity (Roggen et al., 2010) 4 17(+1) 113 $403\pm 63$ $185\pm 48$ $48\pm 10$ $3\pm 2$ $0\pm 0$ non-periodic, complex
SBHAR (Reyes-Ortiz et al., 2016) 30 12(+1) 3 $3\pm 2$ $9\pm 2$ $6\pm 4$ $10\pm 4$ $13\pm 3$ periodic, non-periodic
Wetlab (Scholl et al., 2015) 22 8(+1) 3 $7\pm 5$ $7\pm 3$ $7\pm 2$ $5\pm 2$ $13\pm 2$ complex
WEAR (Bock et al., 2023) 18 18(+1) 12 $1\pm 1$ $1\pm 1$ $1\pm 1$ $1\pm 1$ $30\pm 9$ periodic
Hang-Time (Hoelzemann et al., 2023) 24 5(+1) 3 $154\pm 53$ $22\pm 7$ $10\pm 7$ $3\pm 3$ $3\pm 2$ non-periodic, periodic, complex
RWHAR (Sztyler and Stuckenschmidt, 2016) 15 8 21 $0\pm 0$ $0\pm 0$ $0\pm 0$ $0\pm 0$ $9\pm 1$ periodic
Opportunity
Introduced by Roggen et al. (2010) in 2010, the Opportunity dataset features a sensor-rich recording scenario of four participants performing various activities of daily living. As part of this benchmark analysis we chose to use only the sensor subset associated with the Opportunity challenge Chavarriaga et al. (2013), which consists of 113 sensor axes from IMU and accelerometer body-worn sensors, all sampling at 30Hz. The Opportunity challenge provides two prediction scenarios, i.e. types of annotations, that is modes of locomotion and gestures. For the benchmark analysis presented in this paper, we focus on the latter, which involves predicting 17 types of activities. The Opportunity dataset was recorded under laboratory-like conditions having each participant take part in six different recording sessions. There were two types of sessions: ADL sessions, repeated 5 times, in which participants followed a higher-level protocol and were allowed to interleave activities, and Drill sessions, repeated once, in which participants performed a sequence of activities 20 times. Breaks were included in all recording sessions, resulting in a substantial NULL-class being part of the dataset.
SBHAR
Comprising of 30 participants performing six activities of daily living, the SBHAR dataset augments the base locomotion activities with annotated transitional activities, representing transition periods between two base activities (Reyes-Ortiz et al., 2016). The dataset itself consists of uninterrupted 3D-accelerometer data sampled at 50 Hz, which was captured by a smartphone attached to the waist of each participant. Similar as the Opportunity dataset, the SBHAR dataset was also recorded under laboratory-like conditions, with participants following an experimental protocol that defined repetition count, order and execution length of each activity.
Wetlab
Published by Scholl et al. (2015), the Wetlab dataset provides 3D-accelerometer data from 22 participants. A single inertial sensor, placed in a fixed orientation on the dominant wrist of each participant, captured data at a constant sampling rate of 50Hz. The recording of the dataset took place in a wetlab environment in which each participant performed two DNA extraction experiments. Participants were allowed to take as much time as needed for each experiment. However, following an experimental wetlab protocol, recordings of different participants contain almost identical sequences of consecutive activities. Nevertheless, since not all steps in the experimental protocol were mandatory, some activities were skipped by certain participants. The provided annotated activity streams are two-fold, with one set indicating the steps within the recording protocol and the other representing the recurring base activities, such as peeling, pestling, and cutting. During our benchmark analysis the latter will be used, being the 8 base activities (+ the NULL-class), as it is most comparable to the prediction scenarios present in the other benchmark datasets used.
WEAR
Published by Bock et al. (2023) in 2023, WEAR consists of 3D-accelerometer and egocentric video data from 18 participants performing 18 workout sports activities at 10 outdoor locations. These activities include running-, stretching- and strength-based exercises, with base activities like push-ups being complemented with complex variations that alter and/ or extend the performed movement during the exercise. The 18 activities were divided across multiple recording sessions, with each session consisting of uninterrupted data streams of all modalities, i.e. including a NULL-class. Although participants were assigned a fixed set of workout activities and instructed to perform each activity for approximately 90 seconds, each participant had the freedom to choose the order of activities and take breaks as desired. During each workout session, four inertial sensors all sampling at 50Hz were placed on both left and right wrists and ankles.
Hang-Time
Published by Hoelzemann et al. (2023), Hang-Time features uninterrupted data streams of 24 participants during a basketball practice. The 3D-accelerometer data of each participant was sampled using a smartwatch worn on each participant’s right wrist. Although the dataset provides both locomotion and basketball-related activity labels across three different types of sessions (drills, warm-up and game), we intentionally exclude the locomotion activities during our benchmark analysis, replacing them with a NULL-class label. In the drill sessions, participants followed a strict experimental protocol with predefined activity durations and sequences. In contrast, the warm-up and game sessions allowed participants to commence activities at their discretion in a more real-world recording scenario.
RWHAR
Consisting of 15 participants performing a set of 8 different locomotion activities, the RWHAR dataset (Sztyler and Stuckenschmidt, 2016) was recorded at in-the-wild locations. Unlike the other datasets featured in this benchmark analysis, the RWHAR dataset does not contain a NULL-class, as recordings were trimmed to only contain the activities of relevancy, cutting out the breaks in-between. Except for jumping, each activity was performed the same amount of time, i.e. around 15 minutes. Participants wore a total of 6 body-worn sensors, all set to sample at 50Hz. During this benchmark analysis, we will only be using the 3D-accelerometer and 3D-gyroscope data provided to make modalities comparable to that of the other datasets.
4.2. Training Pipeline
All experiments were initiated employing a Leave-One-Subject-Out cross-validation. This type of validation involves iteratively training on all but one participant’s data and using the hold-out participant’s data during validation until all participants have been evaluated, ensuring that each network architecture is assessed based on its capabilities to generalize across unseen participants. All datasets were windowed using a sliding window of one second with a 50% overlap across windows. For all inertial architectures (Ordóñez and Roggen, 2016; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022) we employed a similar optimization as proposed with the release of the shallow DeepConvLSTM, namely an Adam optimizer paired with a learning rate of $1e^{-4}$, a weight decay of $1e^{-6}$, and Glorot weight initialization (Glorot and Bengio, 2010). To allow each model to converge more properly, we increase the number of epochs to 100 and employ a step-wise learning-rate schedule that multiplies the learning rate by a factor of 0.9 after every 10 epochs. For all architectures, we fixed the hidden dimension of the recurrent layers to employ 128 units and scaled kernel sizes of the convolutional filters according to the relative difference in sampling rate among the different input datasets. In line with how the Attend-and-Discriminate architecture was first introduced, we optimized said architecture using center-loss (Wen et al., 2016) as opposed to a weighted cross-entropy loss, which was used during training of all other inertial architectures. Lastly, as proposed by the authors, we do not shuffle batches during the training of the shallow DeepConvLSTM.
For all three TAL architectures (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) part of this benchmark analysis, we chose to employ hyperparameters reported by the authors that produced best results on the EPIC-KITCHENS dataset (Damen et al., 2022). The hyperparamenters, though optimized for a different modality than inertial data (egocentric videos), have shown to produce competitive results on the WEAR dataset (Bock et al., 2023) and are thus considered a good starting point for evaluating the applicability of the three architectures on other HAR datasets. Nevertheless, given the small size of the tested inertial datasets compared to datasets used by the TAL community, we chose to increase the amount of epochs to 100 throughout all TAL-based experiments. As TAL models are designed to predict both a regression (boundary prediction) and classification target (segment label), loss calculation of the ActionFormer, TemporalMaxer and TriDet models are performed as a weighted combination of a regression loss (IoU loss (Rezatofighi et al., 2019)) and classification loss (focal loss (Li et al., 2020)). The latter, originates from the object detection community, which, like HAR, often times suffers from the background class dominating the class distribution. Furthermore, as we treat each subject’s data as a continuous data stream, we apply the TAL models to predict each subject’s data as a whole. With the input sequences thus being of arbitrary length, the ActionFormer, TemporalMaxer and TriDet model limit their complexity by randomly truncating the input sequence to not exceed a maximum sequence length during training. Furthermore, center sampling is employed when training the three TAL models. This technique makes it such that only time steps within an interval around a segment’s center are to be considered positive, while all other parts are considered negative. Within the TAL community, center sampling has shown to encourage higher scores around activity centers and overall improve results (Tian et al., 2022; Zhang et al., 2020).
As inertial models are tasked to predict on a per-window basis and, in most cases, are trained using shuffled data, the architectures suffer from producing only small coherent segments. Caused by frequently occuring activity switches, the fragmentation ultimately leads to significantly lower mAP scores being produced by the inertial architectures as opposed to the TAL models. Therefore, to remove only short lasting switches, predictions of inertial-based architectures mentioned in this paper were smoothed using a majority-vote filter. The exact size of the majority filter was chosen dataset-specific, being, for example, 15 seconds in the case of the WEAR dataset. Similar to the inertial models, the TAL architectures are tasked to predict class probabilities and segment boundaries of each windowed timestamp. Consequently, without applying any confidence threshold, all predicted activity segments are considered during creation of the prediction timeline causing accuracy of the NULL-class to be significantly low. Therefore, to alleviate this, we apply an optimized confidence threshold on predicted segments of all TAL models. Similarly to the majority filter, the score threshold for each architecture was chosen dataset-specific, being, for instance, 0.2 in the case of the WEAR dataset. This eliminates low scoring segments and substantially lowers the confusion of the NULL-class with the other activities.
5. Results
As part of our experimental evaluation, we provide traditional classification metrics (precision, recall and F1-score), misalignment measures as defined by Ward et al. (2006) and mAP averaged across tIoU thresholds 0.3, 0.4, 0.5, 0.6 and 0.7. All results are the unweighted average across all subjects along the LOSO validation. Experiments were repeated three times employing three different random seeds (1, 2 and 3). Classification metrics are calculated on a per-sample basis as the segmented output of the TAL models and windowed output of the inertial-models need to be translated to a common time granularity. To ensure readability of this work, visualization of the per-class analysis will only include confusion matrices of the SBHAR and RWHAR datasets, as we deemed these two datasets to be the most representative in illustrating the strengths and weaknesses of the TAL architectures when applied to inertial data. Please note that the confusion matrices of the other datasets can be found in the supplementary material. Furthermore, all created plots part of a performed DETAD analysis (Alwassel et al., 2018) on each dataset can be found in our repository.
Per-Dataset Results
Table 2 provides average results of the seven tested architectures across each dataset. One can see that the TAL architectures outperform the inertial architectures across all datasets regarding average mAP. This shows that by being trained to specifically optimize activity boundaries, the different prediction target has resulted in overall more coherent segments which overlap to a larger degree with the ground truth segments. Even though prediction results of the inertial architectures were further smoothed by a majority filter, average mAP is, except for the WEAR dataset, more than halved when compared to that of the TAL architectures. Regarding traditional classification metrics the TAL architectures are able to outperform inertial architectures for four out of six datasets with only the RWHAR (Sztyler and Stuckenschmidt, 2016) and WEAR dataset (Bock et al., 2023) being the exception. These improvements range between $5\%$ for the Opportunity and Wetlab dataset, $10\%$ for the Hang-Time and even as much as 25% in F1-score for the SBHAR dataset. Calculated misalignment ratios show that both inertial and TAL architectures have a similar distribution of errors, with architectures producing better overall classification results also showing overall lower misalignment measures. Only for the RWHAR dataset one can see that TAL architectures show a significantly higher Overfill-Ratio. This might be due to the RWHAR dataset not featuring a NULL-class, which introduces an uncommon prediction scenario for the TAL architectures. Nevertheless, the performed DETAD analysis (Alwassel et al., 2018), which further differentiates amongst the segment-based errors, reveals that inertial architectures suffer more severely from background errors, i.e. confusing activities with the NULL-class. While this effect is decreased for the shallow DeepConvLSTM, TAL architectures show to be able to more reliably differentiate between activities and the NULL-class, and thus more reliably localize activities within the untrimmed sequences.
Table 2. Average LOSO results obtained on six inertial HAR benchmark datasets (Roggen et al., 2010; Reyes-Ortiz et al., 2016; Scholl et al., 2015; Bock et al., 2023; Hoelzemann et al., 2023; Sztyler and Stuckenschmidt, 2016) for four inertial (Ordóñez and Roggen, 2016; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022) and three TAL architectures (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023). The table provides per-sample classification metrics, i.e. Precision (P), Recall (R), F1-Score (F1), misalignment ratios (Ward et al., 2006) and average mAP applied at different tIoU thresholds (0.3:0.1:07). All experiments employed a sliding window of one second with a $50\%$ overlap. Results are averaged across three runs employing different random seeds. The TAL architectures are able to outperform the inertial architectures regarding average mAP on all HAR datasets and result in the highest classification metrics on four out of the six datasets. Best results per dataset are in bold.
Model P ($\uparrow$) R ($\uparrow$) F1 ($\uparrow$) UR ($\downarrow$) OR ($\downarrow$) DR ($\downarrow$) IR ($\downarrow$) FR ($\downarrow$) MR ($\downarrow$) mAP ($\uparrow$)
Opportunity DeepConvLSTM 50.22 33.88 34.41 0.30 17.29 0.78 31.26 0.01 0.23 13.97
Shallow D. 42.08 27.18 26.46 0.24 14.28 0.97 35.06 0.01 0.15 10.61
A-and-D 35.25 48.55 36.35 0.32 15.28 0.45 52.55 0.02 0.35 13.75
TinyHAR 48.09 54.27 47.09 0.34 19.99 0.39 34.01 0.02 0.46 19.78
ActionFormer 54.63 58.78 51.93 0.19 13.34 0.41 33.82 0.01 0.42 51.24
TemporalMaxer 44.44 55.63 44.74 0.21 14.13 0.40 43.78 0.01 0.55 46.31
TriDet 48.72 57.69 48.79 0.23 13.20 0.39 40.07 0.01 0.58 49.70
SBHAR DeepConvLSTM 67.54 63.72 62.31 0.41 7.19 0.59 21.44 0.07 0.10 49.60
Shallow D. 72.98 75.41 71.13 0.60 10.19 0.46 14.23 0.02 0.09 65.15
A-and-D 68.64 71.07 66.49 0.31 9.47 0.45 21.71 0.05 0.11 55.79
TinyHAR 58.91 63.70 56.29 0.31 8.16 0.54 31.67 0.05 0.13 45.38
ActionFormer 87.02 84.37 84.43 0.36 5.41 0.16 5.09 0.00 0.20 95.46
TemporalMaxer 86.52 83.37 83.66 0.41 6.02 0.19 4.41 0.00 0.24 94.39
TriDet 88.95 86.15 86.45 0.38 5.41 0.12 3.95 0.01 0.29 94.75
Wetlab DeepConvLSTM 38.65 47.34 37.87 0.51 7.13 0.69 48.53 0.21 0.64 11.88
Shallow D. 39.01 35.42 34.42 0.51 9.52 1.57 34.51 0.06 0.43 15.40
A-and-D 37.75 55.71 37.49 0.56 9.69 0.63 57.60 0.16 0.57 12.27
TinyHAR 34.31 50.84 31.48 0.61 8.41 1.00 61.26 0.11 0.59 10.05
ActionFormer 40.71 49.25 40.71 0.56 9.41 0.79 52.44 0.07 0.79 33.53
TemporalMaxer 50.43 36.65 37.09 0.59 9.37 0.83 53.59 0.10 0.61 35.72
TriDet 44.13 49.15 42.85 0.58 8.85 0.75 48.63 0.10 0.84 34.05
WEAR DeepConvLSTM 80.68 76.25 75.78 0.28 2.35 0.52 6.68 0.11 0.32 61.03
Shallow D. 80.78 78.91 77.71 0.27 3.23 0.50 5.21 0.04 0.43 67.89
A-and-D 82.34 83.29 80.61 0.20 4.03 0.33 7.18 0.09 0.52 64.78
TinyHAR 81.87 84.23 80.56 0.21 4.77 0.27 9.54 0.10 0.55 63.33
ActionFormer 71.88 76.70 72.43 0.22 6.48 0.65 7.12 0.01 2.06 73.80
TemporalMaxer 69.54 72.80 69.52 0.23 6.87 0.83 5.50 0.01 1.50 69.18
TriDet 73.57 77.54 73.18 0.28 4.79 0.64 6.21 0.03 1.64 77.12
Hang-Time DeepConvLSTM 44.13 33.95 35.25 0.28 10.60 0.88 20.16 0.27 1.46 5.44
Shallow D. 37.97 38.19 36.85 0.35 14.00 1.07 36.38 0.21 2.33 5.00
A-and-D 40.54 43.32 40.39 0.35 14.14 0.72 34.81 0.30 1.44 6.73
TinyHAR 37.09 41.13 36.89 0.41 11.59 0.75 42.90 0.43 1.26 4.73
ActionFormer 49.19 57.57 51.23 0.62 11.63 0.51 47.65 0.48 0.64 29.26
TemporalMaxer 45.01 54.56 47.17 0.71 10.97 0.45 52.71 1.13 0.65 27.86
TriDet 49.59 55.14 50.67 0.73 9.85 0.52 48.55 0.69 0.62 29.24
RWHAR DeepConvLSTM 79.05 81.93 77.56 0.65 2.05 1.09 13.63 1.07 0.00 0.11
Shallow D. 88.59 89.01 86.85 0.31 1.14 0.38 6.93 0.98 0.00 0.00
A-and-D 79.46 82.68 78.09 0.66 2.31 1.17 11.28 0.84 0.00 0.04
TinyHAR 83.59 86.25 82.62 0.53 1.14 0.76 11.65 0.89 0.00 0.02
ActionFormer 63.76 67.64 61.24 2.48 11.12 2.06 11.23 0.09 0.00 65.40
TemporalMaxer 63.20 67.60 60.59 2.59 11.95 1.81 13.96 0.36 0.05 50.60
TriDet 69.27 73.04 67.86 1.48 6.88 2.03 10.24 0.23 0.00 69.98
Per-Class Results
Figures 6 and 7 show the confusion matrices of the TAL and inertial architectures obtained on the SBHAR and RWHAR datasets. With the improvements on the SBHAR being the largest across all datasets, one can see that this increase can be attributed to improved performance on transitional, non-periodic classes like sit-to-stand. These activities are mostly recognisable by their context and surrounding activities and are thus particularly challenging to predict for models that do not rely on temporal dependencies spanning multiple seconds. Since the DeepConvLSTM, Attend-and-Discriminate and TinyHAR models are being trained on shuffled training data and have recurrent parts applied on within-window sequences, the three architectures cannot learn to predict these activities based on surrounding context. Among inertial models, only the shallow DeepConvLSTM is able to more reliably predict these context-based classes as it is trained on unshuffeled data and applies a dimensional flip when training the LSTM. Across all datasets part of this benchmark analysis, the RWHAR dataset yields the least performant results for the TAL architectures. We accredit this primarily due to the absence of a NULL-class in the dataset, which introduces a uncommon prediction scenario for the models. Furthermore, the RWHAR consists of the least amount of segments per subject, limiting the amount of training segments which can be used to optimize the TAL models. Results on the RWHAR dataset are nevertheless surprising as the TAL models show confusion among classes which they do not struggle to predict in other datasets (e.g., lying in the SBHAR dataset) as well as classes that are not similar to each other (e.g., lying and jumping). The obtained results raise the question of whether TAL models are primarily suited for being applied on untrimmed sequences, which (1) include breaks and/ or (2) provide a larger amount of segments than the RWHAR dataset.
Nevertheless, apart from the RWHAR dataset, the TAL models deliver the most consistent results across all classes. Even though, in most cases, the individual per-class accuracies are not the highest when compared to results obtained using the inertial architectures, the inertial architectures are frequently not able to predict all classes reliably, with at least one class showing low prediction accuracy. This is especially evident when examining the results obtained on the Hang-Time and Opportunity datasets, where TAL models are capable of correctly predicting challenging non-periodic activities such as rebounds, passes, opening door and closing door. Furthermore, as also seen in the DETAD analysis, TAL models are overall more capable of differentiating activities from the NULL-class, showing the highest NULL-class accuracy across all datasets. To summarize, the TAL architectures are more reliable in recognizing any kind of actions within an untrimmed sequence, and are less prone to predict fragmented prediction streams or non-existent breaks.
Figure 6. Confusion matrices of the ActionFormer (Zhang et al., 2022b), TemporalMaxer (Tang et al., 2023) and TriDet model (Shi et al., 2023) applied on the SBHAR (Reyes-Ortiz et al., 2016) (top) and RWHAR dataset (Sztyler and Stuckenschmidt, 2016) (bottom) with a one second sliding window and 50% overlap. Results are postprocessed using a score threshold of 0.3 in case of the SBHAR dataset. Note that confusions which are 0 are omitted.
Figure 7. Confusion matrices of the DeepConvLSTM (Ordóñez and Roggen, 2016), shallow DeepConvLSTM (Bock et al., 2021), Attend-and-Discriminate (Abedin et al., 2021) and TinyHAR model (Zhou et al., 2022) applied on the SBHAR (Reyes-Ortiz et al., 2016) (top) and RWHAR dataset (Sztyler and Stuckenschmidt, 2016) (bottom) with a one second sliding window and 50% overlap. Results are postprocessed using a majority filter 5 seconds in case of the SBHAR dataset. Note that confusions which are 0 are omitted.
6. Discussion & Conclusion
This article demonstrated the applicability of vision-based, single-stage Temporal Action Localization for inertial-based Human Activity Recognition. We showed that three state-of-the-art TAL models (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) can be applied in a plain fashion to raw inertial data and achieve competitive results on six popular inertial HAR datasets (Roggen et al., 2010; Reyes-Ortiz et al., 2016; Scholl et al., 2015; Sztyler and Stuckenschmidt, 2016; Hoelzemann et al., 2023; Bock et al., 2023), outperforming in most cases popular models from the inertial community by a significant margin. A previously unexplored metric in inertial-based HAR, mean Average Precision (mAP), reveals that TAL models predict less fragmented timelines compared to inertial models and overall achieve larger degrees of overlap with ground truth segments. Furthermore, TAL models show to predict even non-periodic and complex activities more reliably than inertial architectures, providing consistent results across all types of classes across all datasets. Being one of the key challenges in HAR (Bulling et al., 2014), the TAL architectures further show to be less affected by the unbalanced nature of HAR datasets due to a large NULL-class. Across the five benchmark datasets which offered a NULL-class, the TAL architectures showed to deliver the highest NULL-class accuracy.
Within the past years inertial-based HAR has been working on improving upon methods to model temporal relationships, such as using Transformers as opposed to LSTMs to handle longer sequences (Zhou et al., 2022). Nevertheless, as also evident from this benchmark analysis, the way these mechanisms are used in popular inertial models does not give them the capability to learn temporal context beyond the current timestamp which is to be predicted (Pellatt and Roggen, 2020). This causes predicted timelines to suffer from frequent activity switches. Even though researchers have suggested new metrics (Ward et al., 2011), traditional classification metrics remain the focus of analyses within the inertial community. These metrics, though providing a good intuition about the overall recognition performance, do not provide any insight about the degree of fragmentation of the predicted timeline. With researchers such as Pellatt and Roggen (2020) and Guan and Plötz (2017) having worked on implementing improved training strategies, inertial architectures by default still lack to incorporate context-aware information into their models. Nevertheless, results obtained using the shallow DeepConvLSTM demonstrate that inertial models can be set to learn temporal context, yet require more segment-based optimization to be comparable to TAL architectures as far as average mAP scores.
To summarize, the way TAL architectures are trained does not allow them to be applied in an online fashion. Furthermore, comparing learnable parameters of TAL and inertial models, the former is substantially larger and thus is not suited to be deployed on edge devices. Nevertheless, the retrospective analysis of a fully-captured timeline performed by the TAL community along with fast training times even on large sequences offers an interesting new perspective to inertial-based HAR. Results show that these types of models are leveraging context-based information and are able to outperform popular inertial models on a variety of datasets. Even though models of the two independent communities are sharing similarities, the TAL community offers many unexplored design choices and training concepts, which should be considered to be investigated by the inertial-based HAR community.